OpenStudy (anonymous):

A teacher spends about 6 hours a day teaching in the classroom. What fraction of a day is this?

OpenStudy (anonymous):

1 forth

OpenStudy (anonymous):

one quater

OpenStudy (anonymous):

:)

Parth (parthkohli):

Total hours in a day = 24 Hours taken by the teacher = 6 \(\Large \color{purple}{\rightarrow {6 \over 24} = {1 \over 4} }\)

OpenStudy (aravindg):

6/24=1/4

OpenStudy (lgbasallote):

\[\frac{\text{no. of hours}}{\text{total hours in a day}}\]

OpenStudy (aravindg):

ya divide no of hours with 24 hours

OpenStudy (anonymous):

 HOPE THAT HELPED XD